OpenStudy (anonymous):

what is y squared subtract 10y equal negative 9

OpenStudy (anonymous):

\[^{y2}-10y =-9\]

OpenStudy (anonymous):

y=-1,y=9

OpenStudy (anonymous):

can you show me how to break it down

OpenStudy (radar):

Dipin probably factored the expression by first rearranging the expression to read: y^2-10y+9=0 (adding 9 to both sides) (y-9) (y-1)=0 (factoring) Now if these products equal zero, the one or both of the products equal zero. x-9=0 then x=9 x-1=0 then x=1 It appears that dipin may have lost track of his signs.

OpenStudy (anonymous):

thnx radar, i am sorry for the mistake. y=1,9

OpenStudy (anonymous):

thanks

OpenStudy (radar):

No problem dipin, tavon I hope my changing the name of the variable from y to x did not confuse you, my bad.